OpenStudy (anonymous):

(Medal + Fan. please please I gotta get this work doneee D: and I'll give you a cookie) 1. Choose the correct solution and graph for the inequality. x- 3 < (less than or equal to) 12 a. x< -15 b. x < -4 c. x < 15 d. x<-9

OpenStudy (anonymous):

pretty please please

OpenStudy (anonymous):

All the options satisfy the expression given

OpenStudy (anonymous):

wait huhhh

OpenStudy (shubhamsrg):

Just add 3 to both sides.

OpenStudy (anonymous):

yeah. it would be x=36 right? but where is the answer in it?

OpenStudy (shubhamsrg):

I am not too sure how you got x = 36 ? o.O x-3 < 12 Just add 3 to both sides.

OpenStudy (anonymous):

oh. lmfao. maybe I did it wrong.

OpenStudy (anonymous):

yeah, I defiantly did it wrong.

OpenStudy (anonymous):

C

OpenStudy (anonymous):

C? it's C?

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

okkkkkkkkkkkkkkkkkkkkaaaay doki . ^_^ thank you

OpenStudy (anonymous):

NO problem

OpenStudy (anonymous):

i have nothing else to do for 80 minutes in school for art

OpenStudy (anonymous):

oh wow. I have like this math to do >.<

OpenStudy (anonymous):

i just got done with math what math are you in

OpenStudy (anonymous):

algeba 1

OpenStudy (anonymous):

me 2

OpenStudy (anonymous):

yay! how did chu do in that class ?lol

OpenStudy (anonymous):

math my math teacher is very mean he will yell at you if you dont do your work or dont know how to do the question

OpenStudy (anonymous):

my grade is a 100% because of that

OpenStudy (anonymous):

oh wow lol. could I ask another questionnn?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

\[2. y - 6 \le 2\] a. \[y \le 8\] b. \[y \le -8\] c. \[y \le -8\] d. \[y \le -12\]

OpenStudy (anonymous):

A

OpenStudy (anonymous):

why would it be A? if you allow me asking lol

OpenStudy (anonymous):

just add 6

OpenStudy (anonymous):

ohh. I see.. could I ask another one?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

3. \[\frac{ y }{ 5} > 5 \] a. y > 25 b. y > 5 c. y< 0 d. y < 1

OpenStudy (anonymous):

a times 5 with 5 the faction 5 gets cancel out when you do 5 time 5